The Carbondale Hospital is considering the purchase of a new ambulance. The decision will rest partly on the anticipated mileage to be driven next year. The miles driven during the past 5 years are as follows: 

Year12345
Mileage3.0504,0003,5003.7503.700

a) Using a 2-year moving average, the forecast for year 6 = 3,725 miles (round your response to the nearest whole number) 

b) If a 2-year moving average is used to make the forecast, the MAD based on this = 33.3 miles (round your response to one decimal place). (Hint: You will have only 3 years of matched data) 

c) The forecast for year 6 using a weighted 2-year moving average with weights of 0.35 and 0.65 (the weight of 0.65 is for the most recent period) + 3717.5 miles (round your response to the nearest whole number). 

The MAD for the forecast developed using a weighted 2-year moving average with weights of 0.35 and 0.65  = _______  miles.
Answer #1

Forecast = Sigma(weight for period N * Demand for period N) / Sigma(weights)

Where the highest weights are multiplied with the most recent demand value.

FORECAST 3 = (0.35 * 3050) + 0.65 * 4000) / (0.35 + 0.65) = 3668

FORECAST 4 = (0.35 * 4000) + 0.65 * 3500) / (0.35 + 0.65) = 3675

FORECAST 5 = (0.35 * 3500) + 0.65 * 3750) / (0.35 + 0.65) = 3663

FORECAST 6 = (0.35 * 3750) + 0.65 * 3700) / (0.35 + 0.65) = 3718


MAD


MAD = SIGMA(ABSOLUTE DEVIATION) / N

N = 3

PERIOD

ACTUAL

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

1

3050

2

4000

3

3500

3668

-168

168

4

3750

3675

75

75

5

3700

3663

37

37

SIGMA

280

MAD = 280 / 3 = 93.3